OpenStudy (mtalhahassan2):

Evaluate to the nearest thousandth. a) 3^2x+5 = 5^6-x

OpenStudy (mtalhahassan2):

can someone help me with it

OpenStudy (mtalhahassan2):

plz

OpenStudy (mtalhahassan2):

@peachpi

OpenStudy (anonymous):

This? \[3^{2x+5}=5^{6-x}\]

OpenStudy (mtalhahassan2):

yes

OpenStudy (anonymous):

take the natural log (ln) of both sides. Then apply this rule\[\ln a^b=b \ln a\]

OpenStudy (mtalhahassan2):

what u mean by nature log sorry i forget

OpenStudy (anonymous):

ln

OpenStudy (mtalhahassan2):

can you plz show me a bit more

OpenStudy (anonymous):

this should help you out. http://bosscade.com/c?r=CKUEDE4sd

OpenStudy (mtalhahassan2):

what is that?

OpenStudy (unklerhaukus):

\[\qquad 3^{2x+5}=5^{6-x}\\ \ln(3^{2x+5})=\ln(5^{6-x})\\\qquad\qquad =\]

OpenStudy (unklerhaukus):

now use the rule above

OpenStudy (mtalhahassan2):

ok

OpenStudy (mtalhahassan2):